# Quantum geometric formulation of Brans-Dicke theory for Bianchi I spacetime

### v2

@@ Abstract
We consider a formulation of the Brans-Dicke theory in Jordan’s frame for Bianchi-I spacetime within the framework of loop quantum gravity. The robustness of singularity resolutions due to the quantum effects is explicitly verified in the context of two quantization schemes typically used in the literature. We present an exploration of the effects of quantum geometry on the background dynamics, which is also illustrated through some explicit numerical examples and showing the absence of the singularity.
@@ I Introduction
Loop quantum gravity (LQG) is a candidate of quantum gravity theory, which takes the premise of gravity as a manifestation of geometry of spacetime and systematically constructs a theory of quantum Riemanian geometry with rigor (see, e.g., Ref. [1] for a recent review). LQG stands out as a non-perturbative background independent approach to quantize gravity [2]. At its depth, this theory brings out a fundamental discreteness at Planck scale wherein the underlying geometric observables, such as areas of physical surfaces and volumes of physical regions, are discrete in nature [3, 4, 5, 6]. At present, the studies in cosmology and black holes provide ones of the major avenues for applying and testing the ideas of the theory. The absence of the possibility to design a table top experiment, other than thought experiment, can be easily understood from the energy scale involved in LQG, which is beyond the reach of present day technology. However, it is feasible to apply LQG against observational physics in the context of cosmology, which serves a great purpose here. In fact, the early Universe is a great laboratory for this purpose, so it is relevant to investigate this epoch in this framework.
@@
Loop quantum cosmology (LQC) is an application of LQG techniques to the symmetry reduced spacetime and for an homogeneous spacetime in particular [7]. In LQC the Big Bang singularity is resolved in the sense that it is replaced by a quantum bounce, across which physical macroscopic observables, such as energy density and curvature, which diverge at the Big Bang singularity in classical general relativity (GR), now all remain finite. This non-divergent behavior owes to the fact that the effective quantities in this scenario depends on the fundamental discrete parameters of the theory, in particular, on the fundamental area gap, whose smallest eigenvalue is nonzero [8, 9, 10]. LQC produces a contracting Friedmann-Lemaître-Robertson-Walker (FLRW) universe that bounces back to an expanding one, thus, avoiding the occurrence of a singularity. This is achieved without adding any nontrivial matter component, unlike in the case of classical or matter bounces (see, e.g., Refs. [11, 12]). This quantum bounce occurs purely due to quantum geometric effects, a novel repulsive effective force that is manifest in the quantum corrected Friedmann and Raychaudhuri equations. Also, it is to be noted that in all the different classes of spacetimes permitting different sets of symmetries [13], including the Bianchi and Gowdy models, the singularity is resolved in the framework of LQC [14, 15, 16, 17, 18, 19] (for reviews, see also Refs. [20, 21, 22, 23, 24]).
@@
An important issue that appears in the bouncing scenario is the possible instability in the growth of anisotropic density during the contracting phase. During contraction, the contribution from anisotropic stresses to the Friedmann equation grows much faster than the energy density of the usual fields, like radiation, baryonic matter and cold dark matter. Once the ratio of the anisotropic stresses to the total energy density becomes comparable to one, it is not possible any longer to assume that the background spacetime is approximately isotropic, and this ratio can even become larger than unity, eventually meeting an anisotropic collapse. When considering inhomogeneities, the situation is even more sensitive, since the ratio of the anisotropic stresses to the total energy density becoming larger than unity in this context can indicate the onset of the conjectured Belinski-Khalatnikov-Lifshitz (BKL) chaotic instability [25]. As it is well known, due to the high sensitivity of the dynamics to the initial conditions (neighboring points following very different dynamics), this scenario can lead to the loss of predictivity. In this case, the large anisotropies impact the density fluctuations in a highly inhomogeneous way, spoiling the prediction of a nearly isotropic cosmic microwave background (CMB) [26, 27]. This can be avoided, for example, in some situations when the nearly scale-invariant perturbations are generated after the bounce [28]. Furthermore, in the special case of the Bianchi type I model [29], it was shown that isotropization is an attractor, which makes this scenario a natural choice to investigate viable candidates. This is especially promising in the case of LQC, where we can analyze the Bianchi I background in a bouncing scenario driven by modifications to the Einstein equations in the high energy regime. Great efforts have been devoted towards understanding the quantum corrected dynamics of anisotropic spacetime starting with the Einstein-Hilbert action (see, for example, the Refs. [30, 31, 32, 33, 34, 35, 36]).
@@
A natural question to ask is what it will happen when one goes beyond the Einstein-Hilbert action. One of the salient features of this theory is all the geometrical quantities, such as distance, area and volume can be constructed from the knowledge of the metric. The metric and, hence, the geometry of the spacetime, being a dynamical quantity determined by the distribution of matter and energy. These fundamental properties of the theory remain intact beyond the minimally coupled Einstein-Hilbert action. This motivates us to consider an action with a non-minimal coupling term and to find the corresponding quantum corrected dynamics. As an example toward this goal, we consider the Brans-Dicke theory (BDT). In particular, the techniques of LQG are not limited to a particular action. Based on the loop quantization, some aspects of its implementation in the context of the BDT has been explored for the FLRW spacetime [37, 38, 39], and the presence of a quantum bounce is shown in both the Jordan and Einstein frames. Particularly, an anisotropic Bianchi-I spacetime was considered in LQC in Ref. [40], where a bounce dynamics was shown to emerge. One of the prime focus of the present paper is to formulate the quantum corrected effective dynamics of BDT in homogeneous and anisotropic Bianchi-I spacetime in the Jordan frame, whose study is still missing in the literature, as far as we are aware of. This is a crucial step towards establishing and checking the robustness of the major results of loop quantization in spacetimes with different symmetries and different theories of gravity. In particular, we aim to ultimately check the robustness of the singularity resolution and the consistency of the effective dynamics.
This paper is organized as follows. In Sec. II, we review the Hamiltonian formulation of BDT and present the complete constraint analysis and the associated Lie algebra. Following this, we cast the theory in connection formulation, the Asthekar variables, which is suitable for employing the loop quantization program. Our main results starts from Sec. III, where we present the symmetry reduction of the constraints. As one of our prime focus is to formulate the quantum corrected dynamics of BDT in Bianchi-I spacetime. In Sec. IV we project the full BDT dynamics in Bianchi-I spacetime at the classical level before quantization. The classical dynamics is found by taking the Poisson flow of the phase-space variables with the scalar Hamiltonian constraint. In Sec. V, we find the quantum corrected dynamics using two alternative and complementary $\bar{\mu}$ quantization schemes which have been considered in the literature in other contexts. In Sec. VI, we discuss the dynamics of the Brans-Dicke theory for Bianchi LQC models in the two quantization schemes considered in this paper. Our final considerations and conclusions are presented in Sec. VII. Throughout this work, we will adopt the convention $M_{Pl}=1$, where $M_{Pl}=\frac{1}{8\pi G}$ is the reduced Planck mass.
II Hamiltonian formulation of Brans-Dicke theory and constraint analysis
The action for the Brans-Dicke theory is given by
@@ II Hamiltonian formulation of Brans-Dicke theory and constraint analysis
where $R$ is the Ricci scalar, $g_{\mu\nu}$ is the metric and $g$ is its determinant, the scalar field $\phi$ is the Brans-Dicke (BD) scalar and $\omega$ is a coupling constant. The equations of motion are obtained by varying the action, Eq. (1), independently, with respect to the metric $g_{\mu\nu}$ and the scalar field $\phi$, which, in natural units, give, respectively,
$\phi G_{\mu\nu}$ $=$ $\gradient_{\mu}\gradient_{\nu}\phi-g_{\mu\nu}\Box\phi+$ (2)
$+$ $\frac{\omega}{\phi}\left[\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}g_{\mu\nu}(\gradient\phi)^{2}\right],$
$R$ $+$ $\frac{2\omega}{\phi}\Box{\phi}-\frac{w}{\phi^{2}}(\partial_{\mu}\phi)\partial^{\mu}\phi=0,$ (3)
where $R_{\mu\nu}$ is the Ricci tensor and $G_{\mu\nu}\equiv R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R$ is the Einstein tensor. However, the present format of covariant formulations of the theory is not suitable for loop quantization. In the Lagrangian formulation, space and time are treated on equal footing, but the construction of the canonical formula demands splitting of spacetime into space and time. It is the Arnowitt-Deser-Misner (ADM) formalism [41, 42, 43] with the $3+1$ decomposition that comes to rescue the description in terms of the first-order Hamilton’s equation of the phase space variables. In brief, this can be achieved by decomposing the spacetime manifold endowed with a metric, designated by the pair $(\mathcal{M},g_{\mu\nu})$, into $(\Sigma$$,\mathbb{R})$, where $\Sigma$ represents the spatial three-dimensional hypersurface and $\mathbb{R}$ stands for the real valued time. Algebraically, it is represented by splitting the metric as
$q_{\mu\nu}=g_{\mu\nu}-n_{\mu}n_{\nu},$ (4)
where $q_{\mu\nu}$ is the intrinsic spatial metric that defines the hypersurface $\Sigma$ and $n_{\mu}$ is the normal vector to the hypersurface, with the property $n_{\mu}n^{\nu}=-1$.
A careful observation of the splitting of the metric into its intrinsic spatial metric $q_{\mu\nu}$ and the normal vector $n_{\mu}$ to the hypersurface naturally leads to two notions of curvature [44]: the intrinsic curvature, $\mathcal{{R}}$, purely defined in terms of commutator of covariant derivatives strictly sitting on the three-dimensional hypersurface, and the extrinsic curvature, K, defined in terms of the covariant derivative of the normal vector to the hypersurface. Due to the Gauss-Codazzi equation [44], a relation between the two is given in an elegant way,
$R$ $=$ $K_{ab}K^{ab}-K^{2}+\mathcal{R}+\frac{2}{\sqrt{-g}}\partial_{\mu}(\sqrt{-g}n^{\mu}K)$ (5)
-- $\frac{2}{N\sqrt{q}}\partial_{a}(\sqrt{q}q^{ab}\partial_{b}N).$
The set of corresponding conjugate momenta, in terms of the newly defined geometrical objects, for the BDT action given by Eq. (1), are
@@
$\pi\equiv\frac{\delta\mathcal{L}}{\delta\dot{\phi}}=-\sqrt{q}\left[K-\frac{w}{N\phi}(\dot{\phi}-N^{c}\partial_{c}\phi)\right],$ (7)
$p_{N}(x)\equiv\frac{\delta\mathcal{L}}{\delta N(x)}=0,$ (8)
$p_{a}(x)\equiv\frac{\delta\mathcal{L}}{\delta\dot{N}^{a}(x)}=0.$ (9)
Thus, given the Lagrangian, the Hamiltonian can be obtained by the Legendre transformation,
$H_{\rm total}$ $=$ $\int_{\Sigma}d^{3}x\sqrt{q}(\dot{q}_{ab}p^{ab}+\dot{\phi}\pi-\mathcal{L}(q_{ab},\dot{q}_{ab})),$
@@
$H_{\rm total}$ $=$ $\int_{\Sigma}d^{3}x\,N^{a}\left[-2D^{b}p_{ab}+\pi\partial_{a}\phi\right]$
$+$ $\int_{\Sigma}d^{3}x\,N\left\{\frac{2}{\sqrt{q}\phi}\left[p_{ab}p^{ab}-\frac{1}{2}p^{2}+\frac{(p-\pi\phi)^{2}}{2(3+2\omega)}\right]\right.$
$+$ $\left.\frac{\sqrt{q}}{2}\left[-\phi R+\frac{w}{\phi}(D_{a}\phi)D^{a}\phi+2D_{a}D^{a}\phi\right]\right\}.$
In terms of the Asthekar variables, with the densetized triads $E^{a}_{i}\equiv\sqrt{e}e^{a}_{i}$, curvature tensor $F^{i}_{ab}\equiv\partial_{[a}A^{i}_{b]}+\epsilon^{i}_{kl}A^{k}_{l}A^{k}_{b}$, with the Yang-Mills connection $A^{i}_{a}\equiv\Gamma^{i}_{a}+\gamma K^{i}_{a}$, the constraints of the BDT in the Yang-Mills phase space read
@@
$\mathcal{V}_{a}$ $=$ $\frac{1}{\gamma}F^{i}_{ab}E^{b}_{i}\approx 0,$ (13)
@@
$+$ $\frac{1}{3+2\omega}\left[\frac{(\tilde{K}^{i}_{a}E^{a}_{i})^{2}}{\phi\sqrt{q}}+\frac{\tilde{K}^{i}_{a}E^{a}_{i}}{\sqrt{q}}\Pi+\frac{\Pi^{2}\phi}{\sqrt{q}}\right]\approx 0,$
which are the Gauss, diffeomorphism and scalar constraints, respectively. In the above expressions, $\gamma$ is the Barbero-Immirzi parameter and $\tilde{K}^{i}_{a}=e^{b}_{i}\tilde{K}K_{ab}$ with
@@
III Symmetry Reduction of BDT for the Bianchi I model

### v3

@@ Abstract
This paper investigates Bianchi I spacetimes within the Jordan frame of Brans-Dicke theory, incorporating the framework of effective loop quantum gravity. After developing general formulas, we analyze the robustness of classical singularity resolution due to quantum geometric effects using two common quantization schemes. We then compare the resulting physical properties. We find that both schemes replace classical singularities with regular quantum bounces. Notably, in contrast to similar studies based on general relativity, we find that all three directional scale factors of the Bianchi I spacetimes increase and after the quantum bounce they reach values similar to their initial values, leading to a merging with classical spacetimes in both schemes.
@@ I Introduction
Loop quantum gravity (LQG) is a candidate of quantum gravity theory, which takes the premise of gravity as a manifestation of geometry of spacetime and systematically constructs a theory of quantum Riemanian geometry with rigor (see, e.g., Ref. [1] for a recent review). LQG stands out as a nonperturbative background independent approach to quantize gravity [2, 3, 4, 5, 6]. At its depth, this theory brings out a fundamental discreteness at Planck scale wherein the underlying geometric observables, such as areas of physical surfaces and volumes of physical regions, are discrete in nature [7, 8, 9, 10]. At present, the studies in cosmology and black holes provide ones of the major avenues for applying and testing the ideas of the theory. The absence of the possibility to design a table top experiment, other than thought experiment, can be easily understood from the energy scale involved in LQG, which is by far beyond the reach of present technology. However, it is feasible to apply LQG against observational physics in the context of quantum cosmology. In fact, the early Universe provides a natural laboratory for this purpose, and it is very interesting and important to investigate the Planckian physics of the very early Universe in the framework of LQG.
@@
Loop quantum cosmology (LQC) is an application of LQG techniques to the symmetry reduced spacetimes  [11, 12, 13, 14, 15, 16]. In LQC the big bang singularity is resolved in the sense that it is replaced by a quantum bounce, across which physical observables, such as energy density and curvatures, which diverge at the big bang singularity in classical general relativity (GR), now all remain finite. This nondivergent behavior owes to the fact that in LQG the area operator has a discrete spectrum of eigenvalues with a nonzero area gap [7, 8, 9, 10]. It is this nonzero gap that prevents the formation of spacetime singularities [17, 18, 19]. This is quite similar to the simple harmonic oscillator, whose ground state energy is not zero quantum mechanically, although classically it does. Remarkably, LQC produces a contracting Friedmann-Lemaître-Robertson-Walker (FLRW) Universe that bounces back to an expanding one, thus, avoiding the occurrence of a singularity. This is achieved without adding any nontrivial matter component, unlike in the case of classical or matter bounces (see, e.g., Refs. [20, 21, 22]). This quantum bounce occurs purely due to quantum geometric effects, a novel repulsive effective force that is manifest in the quantum corrected Friedmann and Raychaudhuri equations. The massless scalar field that acts as a clock in this setting, is identified with the inflaton that drives a nearly exponential expansion of the Universe given by the effective dynamics. The preinflationary background dynamics have been extensively studied in detail for various forms of potentials in the Refs. [23, 24, 25, 26, 27]. In addition, in all spacetimes studied so far, which permit different symmetries, including the Bianchi and Gowdy models [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], the singularity is resolved [11, 12, 13, 14, 15, 16].
@@
An important issue that appears in the bouncing scenario is the possible instability in the growth of anisotropic density during the contracting phase. As a matter of fact, during this contraction phase, the contribution from anisotropic stresses to the Friedmann equation grows much faster than the energy density of the usual fields, like radiation, baryonic matter and cold dark matter. Once the ratio of the anisotropic stresses to the total energy density becomes comparable, it is not possible any longer to assume that the background spacetime is approximately isotropic, and this ratio can even become larger than unity, eventually leading to an anisotropic collapse. When considering inhomogeneities, the situation is even more sensitive. In particular, when the ratio of the anisotropic stresses to the total energy density is larger than unity, it indicates the onset of the conjectured Belinski-Khalatnikov-Lifshitz chaotic instability [48]. As is well-known, due to the high sensitivity of the dynamics to the initial conditions (neighboring points following very different dynamics), this scenario can lead to the loss of predictivity. In this case, the large anisotropies impact the density fluctuations in a highly inhomogeneous way, spoiling the prediction of a nearly isotropic cosmic microwave background [49, 50]. This can be avoided, for example, in some situations when the nearly scale-invariant perturbations are generated after the bounce [51]. Furthermore, in the special case of the Bianchi type I model [52], it was shown that isotropization is an attractor, which makes this scenario a natural choice to investigate viable candidates. This is especially promising in the framework of LQC, where we can analyze the Bianchi I background in a bouncing scenario driven by modifications to the Einstein equations in the high energy regime. Great efforts have been devoted towards understanding the quantum corrected dynamics of anisotropic spacetime starting with the Einstein-Hilbert action (see, for example, Refs. [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47] and references therein).
@@
A natural question to ask is what will happen when one goes beyond the Einstein-Hilbert action. One of the salient features of GR is that all the geometrical quantities, such as distance, area and volume can be constructed from the knowledge of the metric. The metric and, hence, the geometry of the spacetime, being a dynamical quantity that is determined by the distribution of matter and energy. These fundamental properties of the theory remain intact even beyond the minimally coupled Einstein-Hilbert action. This motivates us to consider an action with a nonminimal coupling term and to find the corresponding quantum corrected dynamics. As an example toward this goal, we consider the Brans-Dicke theory (BDT). In particular, the techniques of LQG are not limited to a particular action. Based on the loop quantization, some aspects of its implementation in the context of the BDT has been explored for the FLRW spacetime [53, 54, 55, 56, 57] and the presence of a quantum bounce is shown in both the Jordan and Einstein frames.
One of the prime focus of the present paper is to formulate the quantum corrected effective dynamics of BDT in homogeneous and anisotropic Bianchi I spacetimes in the Jordan frame, whose study is still missing in the literature, as far as we are aware of. This is a crucial step towards establishing and checking the robustness of the major results of loop quantization in spacetimes with different symmetries and different theories of gravity. In particular, we aim to ultimately check the robustness of the singularity resolution and the consistency of the effective dynamics.
This paper is organized as follows. In Sec. II, we briefly review the Hamiltonian formulation of BDT and provide the necessary formulas for the studies of Bianchi I spacetimes in Jordan’s frame of the theory. In Sec. III, we present the symmetry reduction of the classical Hamiltonian constraint for Bianchi I spacetimes and write down the corresponding classical dynamical equations. In Sec. IV, we study the quantum corrected dynamics using two alternative and complementary quantization schemes, which have been considered in the literature, and write down the dynamical equations for each of the schemes. In order to study quantum effects of matter fields, in the effective Hamiltonian, we also include a massless scalar field $\Phi$, in addition to the Brans-Dicke (BD) scalar field $\phi$. In Sec. V, we study numerically the dynamics of the BDT for Bianchi LQC models in both quantization schemes and show explicitly that the classical spacetime singularity is indeed resolved in both schemes and replaced by a quantum bounce. Our main conclusions are presented in Sec. VI. Throughout this work, we will adopt the convention $M_{\rm Pl}=1$ (unless in the cases where possible confusions can raise), where $M_{\rm Pl}\equiv 1/\sqrt{8\pi G}$ is the reduced Planck mass.
II Hamiltonian formulation of Brans-Dicke theory
To facilitate our investigations to be carried out in the next sections, in this section we shall provide a very brief review of BDT and for more details, we refer the readers to [53, 54, 55, 56, 57]. The BDT is described by the action
@@ II Hamiltonian formulation of Brans-Dicke theory
where $R$ is the Ricci scalar, $g_{\mu\nu}$ the metric and $g$ its determinant. The BD scalar field is denoted by $\phi$ and $\omega$ is the coupling constant, which is highly constrained observationally [58]. The present format of covariant formulations of the theory is not suitable for loop quantization. In the Lagrangian formulation, space and time are treated on equal footing, but the construction of the canonical formula demands splitting of spacetime into space and time. It is the Arnowitt-Deser-Misner formalism [59, 4, 60] with the $3+1$ decomposition that comes to rescue the situation by providing a description in terms of the first-order Hamilton’s equations of the phase space variables. Formally, this can be represented by decomposing the spacetime manifold, $\mathcal{M}$ endowed with a metric, designated by the pair $(\mathcal{M},g_{\mu\nu})$, into $(\Sigma$$,\mathbb{R})$, where $\Sigma$ represents the spatial three-dimensional hypersurface and $\mathbb{R}$ stands for the real valued time.
The splitting of the metric into its intrinsic spatial metric $q_{ab}\;(a,b=1,2,3)$ and the normal vector $n_{\mu}\;(n_{\nu}n^{\nu}=-1)$ to the hypersurface naturally introduces the notion of extrinsic curvature, $K_{ab}$, defined in terms of covariant derivative of the normal vector to the hypersurface as projected to the hypersurface [61]. Let us denote q as the determinat of $q_{ab}$ and $N$ denotes the lapse function. Then, the corresponding conjugate momenta for the BDT action given by Eq. (1) are
@@
$p_{\phi}\equiv\frac{\delta\mathcal{L}}{\delta\dot{\phi}}=-\sqrt{q}\left[K-\frac{w}{N\phi}(\dot{\phi}-N^{c}\partial_{c}\phi)\right],$ (3)
$p_{N}(x)\equiv\frac{\delta\mathcal{L}}{\delta\dot{N}(x)}=0,$ (4)
$p_{a}(x)\equiv\frac{\delta\mathcal{L}}{\delta\dot{N}^{a}(x)}=0,$ (5)
where the Lagrangian density $\mathcal{L}$ is given by [62]
$\mathcal{L}$ $=$ $\frac{\sqrt{q}}{2\kappa^{2}}\Bigg{[}N\phi\left(\mathcal{R}+(K^{ab}K_{ab}-K^{2})\right)+2q^{ab}D_{a}ND_{b}{\phi}$ (6)
-- $\frac{\omega}{N\phi}\left(N^{2}q^{ab}D_{a}{\phi}D_{b}\phi-(\dot{\phi}-N^{a}D_{a}{\phi})^{2}\right)$
-- $2K(\dot{\phi}-N^{a}D_{a}\phi)\Bigg{]}.$
Here the trace of the extrinsic curvature tensor is $K\equiv K^{a}_{a}=q^{ab}K_{ab}$ and the spatial covariant derivative is defined as $D_{a}\equiv q^{b}_{a}\nabla_{b}$. For a given Lagrangian, the Hamiltonian can be obtained by the Legendre transformation,
$H_{\rm total}$ $=$ $\int_{\Sigma}d^{3}x\sqrt{q}(\dot{q}_{ab}p^{ab}+\dot{\phi}p_{\phi}-\mathcal{L}(q_{ab},\dot{q}_{ab})),\;\;\;$ (7)
@@
$H_{\rm total}$ $=$ $\int_{\Sigma}d^{3}x\,N^{a}\left[-2D^{b}p_{ab}+p_{\phi}\partial_{a}\phi\right]$
$+$ $\int_{\Sigma}d^{3}x\,N\left\{\frac{2}{\sqrt{q}\phi}\left[p_{ab}p^{ab}-\frac{1}{2}p^{2}+\frac{(p-p_{\phi}\phi)^{2}}{2(3+2\omega)}\right]\right.$
$+$ $\left.\frac{\sqrt{q}}{2}\left[-\phi R+\frac{w}{\phi}(D_{a}\phi)D^{a}\phi+2D_{a}D^{a}\phi\right]\right\},$
In terms of the Asthekar variables, with the densetized triads $E^{a}_{i}\equiv\sqrt{e}e^{a}_{i}$, curvature tensor $F^{i}_{ab}\equiv\partial_{[a}A^{i}_{b]}+\epsilon^{i}_{kl}A^{k}_{l}A^{k}_{b}$, and the Yang-Mills connection $A^{i}_{a}\equiv\Gamma^{i}_{a}+\gamma K^{i}_{a}$, the constraints of the BDT in the Yang-Mills phase space read
@@
$\mathcal{V}_{a}$ $=$ $\frac{1}{\gamma}F^{i}_{ab}E^{b}_{i}+p_{\phi}\partial_{a}\phi\approx 0,$ (10)
@@
$+$ $\frac{1}{3+2\omega}\left[\frac{(\tilde{K}^{i}_{a}E^{a}_{i})^{2}}{\phi\sqrt{q}}+2\frac{\tilde{K}^{i}_{a}E^{a}_{i}}{\sqrt{q}}p_{\phi}+\frac{p_{\phi}^{2}\phi}{\sqrt{q}}\right]$
$+$ $\frac{\omega}{2\phi}\sqrt{q}(D_{a}\phi)D^{a}\phi+\sqrt{q}D_{a}D^{a}\phi\approx 0,$
which are often referred to as the Gauss, diffeomorphism and scalar constraints, respectively. In the above expressions, $\gamma$ is the Barbero-Immirzi parameter and $\tilde{K}^{i}_{a}=e^{b}_{i}\tilde{K}K_{ab}$ with
@@
III Symmetry Reduction of Classical BDT for Bianchi I Spacetimes
The four-dimensional Bianchi I spacetimes described by the metric
$ds^{2}$ $=$ $-N^{2}dt^{2}+q_{ab}dx^{a}dx^{b}$ (14)
$=$ $-N^{2}dt^{2}+\sum_{I=1}^{3}{a_{I}^{2}\left(dx_{I}\right)^{2}},$